OpenStudy (andrea6201):

Find the derivative of f(x) = -6 divided by x at x = 12.

OpenStudy (luffingsails):

Another way to look at: \[-6/x\] is to think of it as \[-6x^{-1}\] That should make the derivative a bit simpler.

OpenStudy (andrea6201):

Well, I am using the difference quotient, so when doing the part involving finding f(x+h), i'm not sure how to do that with the negative exponent. What i'm really having trouble with is after setting the equation up using difference quotient: \[\frac{\frac{ 6 }{ x + h } - (-\frac{ 6 }{ x }) }{ h }\] I don't know what do after this.

OpenStudy (lord_box):

\[\frac{ \frac{ -6 }{ x + h } - \frac{ -6 }{ x }}{ h }\] Careful with your signs

OpenStudy (luffingsails):

Try finding a common denominator. For instance multiple the left side by \[\frac{x}{x}\] and the right side by \[\frac{x+h}{x+h}\] that will allow you to add them and you can begin to simplify.

OpenStudy (andrea6201):

Right, sorry about that, thought I had written the negative in the first fraction :)

OpenStudy (andrea6201):

OK so, \[\frac{ \frac{ -6x ^{2} - 6x - 6h }{ x ^{2}+hx } } { h }?\]

OpenStudy (andrea6201):

Wait a second, I messed up, it would be: \[\frac{ \frac{ -6x + 6x + 6h }{ x ^{2}+hx } }{ h }\] and then i'd cancel out the -6x + 6x to only have 6h on the very top top of the equation yes?

OpenStudy (luffingsails):

Right, and then you would be able to divide by h in each of the terms. What do you have left?

OpenStudy (andrea6201):

What do you mean by divide by h? I was wondering if I could multiply by 1/h.

OpenStudy (luffingsails):

Right, same thing in this case.

OpenStudy (andrea6201):

Ok so would that cancel out the h, in 6h? and i'd be left with 6/x^2 + hx?

OpenStudy (andrea6201):

Wait no I was cross-multiplying. Okay wait, so I would multiply the entire denominator by h yes? And get 6h / hx^2 + h^2x????

OpenStudy (andrea6201):

OH WAIT I GOT IT

OpenStudy (andrea6201):

\[\frac{ 6h }{h(x^2 +hx) }\] cancel out the h's, and i'm left with \[\frac{ 6 }{ x ^{2} +hx}\]

OpenStudy (andrea6201):

Ok and then, I substitute 0 into h, get 6/x^2, replace x with 12 and my simplified answer is 1/24. Would this be correct?

OpenStudy (legomyego180):

(1/24) is correct

OpenStudy (andrea6201):

Yes, thank you very much!

OpenStudy (luffingsails):

Sorry, I couldn't help you to the end but excellent work on your part!